# Deciphering Movement: Unified Trajectory Generation Model for Multi-Agent

### v1

@@ Abstract
Understanding multi-agent behavior is critical across various fields. The conventional approach involves analyzing agent movements through three primary tasks: trajectory prediction, imputation, and spatial-temporal recovery. Considering the unique input formulation and constraint of these tasks, most existing methods are tailored to address only one specific task. However, in real-world applications, these scenarios frequently occur simultaneously. Consequently, methods designed for one task often fail to adapt to others, resulting in performance drops. To overcome this limitation, we propose a Unified Trajectory Generation model, UniTraj, that processes arbitrary trajectories as masked inputs, adaptable to diverse scenarios. Specifically, we introduce a Ghost Spatial Masking (GSM) module embedded within a Transformer encoder for spatial feature extraction. We further extend recent successful State Space Models (SSMs), particularly the Mamba model, into a Bidirectional Temporal Mamba to effectively capture temporal dependencies. Additionally, we incorporate a Bidirectional Temporal Scaled (BTS) module to comprehensively scan trajectories while maintaining the temporal missing relationships within the sequence. We curate and benchmark three practical sports game datasets, Basketball-U, Football-U, and Soccer-U, for evaluation. Extensive experiments demonstrate the superior performance of our model. To the best of our knowledge, this is the first work that addresses this unified problem through a versatile generative framework, thereby enhancing our understanding of multi-agent movement. Our datasets, code, and model weights are available at https://github.com/colorfulfuture/UniTraj-pytorch.
@@ 1 Introduction
Analyzing the behavior of multiple agents is invaluable across various domains, including autonomous driving [19, 32], video surveillance [21, 20], and sports analytics [70, 74]. To decipher agent behaviors, these applications rely heavily on key tasks such as multi-object tracking [44], person re-identification [86], trajectory modeling [51, 51], and action recognition [93, 18]. Among these tasks, trajectory modeling is particularly straightforward and effective for understanding the movements of agents. Despite its inherent challenges, such as the complexity of dynamic environments and the subtle agent interactions, significant advancements have been made recently. These advancements are concentrated in three main areas: trajectory prediction, imputation, and spatial-temporal recovery. As depicted in Figure 1, these tasks differ in their input constraints and requirements.
Significant developments have been made recently in modeling multi-agent trajectories, however, most approaches are specialized for specific tasks. For example, numerous studies [83, 84, 60, 48, 36, 12, 96, 28, 6, 8, 66, 13, 64, 54, 55, 82] have focused on multi-agent trajectory prediction, achieving promising results in public datasets. Nonetheless, these approaches often struggle to generalize to
@@
other trajectory-related tasks, such as trajectory imputation and spatial-temporal recovery. These tasks require exploiting both forward and backward spatial-temporal dependencies, which are not typically addressed in prediction-focused models. While some studies [43, 95] have tackled the multi-agent trajectory imputation problem, they often overlook the future trajectories of agents. This omission limits their practical utility in complete movement understanding, where predicting future trajectories is crucial for downstream planning rather than merely reconstructing historical trajectories. Although some recent efforts [85, 59] have attempted to integrate imputation and prediction, they often employ a multi-task framework that relies solely on masked observations, resulting in a lack of diversity in missing formats. Given that any situation might happen in real practice, it is crucial to develop a general method that can accommodate various scenarios. This raises two pivotal questions: (1) How can we unify these disparate but relevant tasks into a general framework that accommodates different settings? (2) How can we effectively model these trajectories with varying input formulations?
Prompted by these questions, we revisit these tasks and introduce the Unified Trajectory Generation model (UniTraj), which integrates these tasks into a general scheme. Specifically, we merge different input types into a single unified formulation: an arbitrary incomplete trajectory with a mask that indicates the visibility of each agent at each time step. We uniformly process the inputs of each task as masked trajectories, aiming to generate complete trajectories based on the incomplete ones. To model spatial-temporal dependencies across various trajectory formulations, we introduce a Ghost Spatial Masking (GSM) module embedded within a Transformer-based encoder for spatial feature extraction. Leveraging the notable capability of recent popular State Space Models (SSMs), especially the Mamba model, we adapt and enhance it into a Bidirectional Temporal Mamba encoder for long-term multi-agent trajectory generation. Furthermore, we propose a simple yet effective Bidirectional Temporal Scaled (BTS) module that comprehensively scans trajectories while preserving the integrity of temporal relationships within the sequence. Owing to a lack of benchmarks in this novel setting, we benchmark and release three sports game datasets: Basketball-U, Football-U, and Soccer-U, to facilitate evaluation. In summary, the contributions of our work are as follows:
@@
  • We curate and benchmark three sports game datasets, Basketball-U, Football-U, and Soccer-U, and establish strong baselines for this integrated challenge.
@@
2 Related Work
Trajectory Prediction.
Trajectory prediction aims to predict the future movements of agents conditioned on their historical observations. The biggest challenge is modeling the social interactions among agents, which has driven the development of various methods. A classic approach, Social-LSTM [2], introduced a pooling layer to enable information sharing. Subsequent methods [72, 92, 34, 84] have employed similar designs to extract comprehensive interaction features. More recent studies [37, 67, 50, 65, 79, 41, 7] have utilized Graph Neural Networks (GNNs), treating agents as nodes to model social interactions. The inherent uncertainty and diversity of future trajectories have also led to the adoption of generative models, including Generative Adversarial Networks (GANs) [30, 61, 40, 4], Conditional Variational Autoencoders (CVAEs) [47, 80, 35, 62, 81], and Diffusion models [29, 48, 36].
While these models have shown promising results, most assume that observations are complete, overlooking potential missing observations. Recent works [58, 85, 39] have highlighted the issue of missing observations, incorporating both imputation and prediction tasks into a multi-task framework. However, these efforts primarily address visible gaps in historical trajectory data, without considering the more complex scenarios of missing data that occur in practice. Moreover, prediction may not always be the ultimate objective in applications such as similarity analysis or action localization. Our work aims to tackle a broader range of challenges within trajectory modeling and introduces a new benchmark to better address these complexities.
Trajectory Imputation and Spatial-Temporal Recovery.
Imputation is a classic and extensively explored task, with time-series imputation receiving the most attention. Traditional statistical techniques include replacing missing values with the mean or median value [1], as well as methods like linear regression [5], k-nearest neighbors [69, 10], and the expectation-maximization (EM) algorithm [24, 52]. These methods often suffer from limited generalization ability due to their reliance on rigid priors. In response, more flexible frameworks have emerged, employing deep learning techniques for sequential data imputation, such as autoregressive imputation using RNNs [88, 11], or the use of GANs, VAEs, and Diffusion models [87, 45, 46, 58, 49, 68, 78, 14, 89] to generate reconstructed sequences. However, only a few studies have specifically addressed trajectory imputation for multi-agent movements. For instance, AOMI [43] introduces a non-autoregressive imputation method, while GMAT [90] develops a hierarchical model to generate macro-intent labels for sequence generation. Additionally, Graph Imputer [53] has utilized forward and backward information to model the distribution of imputed trajectories in soccer games.
The task of trajectory spatial-temporal recovery, while similar to trajectory imputation in its focus on incomplete trajectories, differs in its broader goal of reconstructing complete spatial-temporal sequences. This task requires not only filling gaps but also mining the intrinsic relations between different agents’ trajectories, which is crucial for real-world applications like arrival time estimation [22, 17] and trajectory similarity computation [31]. A major benchmark involves recovering trajectory data from GPS points [16, 76, 77, 15], while few studies focus on multi-agent contexts.
Recent research has integrated trajectory imputation and prediction tasks. For instance, INAM [58] introduces an imitation learning paradigm that handles both tasks asynchronously, while GC-VRNN [85] proposes a multi-task framework using three different GCN layers and a variational RNN to conduct trajectory imputation and prediction simultaneously. Nevertheless, these studies primarily aim to forecast trajectories based on missing observations, whereas our work aims for a more general goal that is not restricted to specific input formats or tasks.
State Space Models.
The core concept of state space sequence models (SSMs) [26, 27] is to link input and output sequences through a latent state. Recently, a promising variant of SSMs, Mamba architecture [25], which integrates time-varying parameters, has been introduced to enhance the efficiency. This architecture has inspired various recent studies in different computer vision tasks. For instance, in works [42, 57, 73, 9, 97], Mamba has been applied to effectively process image patches, learn visual representations, and achieve impressive performance in downstream tasks, proving it as an effective backbone.
More relevant to our work, several Mamba-based methods have been developed to solve the problem of sequence generation. For instance, Motion Mamba [94] employs SSMs for efficient and long-term motion generation, incorporating multiple isolated SSM modules within a symmetric U-Net architecture. Similarly, Traj-LLM [38] uses the Mamba model to develop a lane-aware probabilistic learning module and integrates Large Language Models (LLMs) for trajectory prediction in autonomous driving. Additionally, Simple-Mamba [75] tokenizes the time points of each variable and uses the Mamba model for time series forecasting. However, the full potential of the Mamba model in trajectory modeling remains underexplored. In this work, we leverage and expand upon the Mamba model to learn temporal dependencies from both directions and introduce a novel Bidirectional Temporal Scaled (BTS) module for a comprehensive trajectory scan.
@@ 3.1 Problem Definition
Table 1: Overall architecture of our UniTraj model. The encoders extract agent features and derive latent variables, while the decoder generates the complete trajectory using the sampled latent variables and agent features.
@@ 3.2 Unified Trajectory Generation Model
where $\varphi_{x}(\cdot)$ and $\varphi_{y}(\cdot)$ are projection functions with weights $\mathbf{W}_{x}\in\mathbb{R}^{8\times D}$ and $\mathbf{W}_{y}\in\mathbb{R}^{8\times D}$, $\odot$ represents element-wise multiplication, and $\oplus$ indicates concatenation. We implemented $\varphi_{x}(\cdot)$ and $\varphi_{y}(\cdot)$ using MLPs. This approach allows us to incorporate location, velocity, visibility, and category information for extracting spatial features for subsequent analyses.
@@
Unlike other sequential modeling tasks, it is crucial to consider dense social interactions, especially in sports games. Existing studies on human interactions predominantly use attention mechanisms like cross-attention and graph attention to capture these dynamics. However, given that we are addressing a unified problem with arbitrary incomplete inputs, it is essential for our model to learn the spatial-temporal missing patterns. While some studies [95, 85] utilize edge graphs with attached labels to extract missing features, this approach is resource-intensive and requires the construction of an additional graph layer. In contrast, we propose a novel and efficient Ghost Spatial Masking (GSM) module to abstract and summarize the spatial structures of missing data. This module can be smoothly integrated into the Transformer without complicating the model structure.
The Transformer [71] was originally proposed to model the temporal dependencies for sequential data, and we adopt the multi-head self-attention design along the spatial dimension. At each time step, we treat each agent’s embedding as a token and input $N$ agent tokens into the Transformer encoder. To address various missing scenarios at each time step, we repeat the mask to the feature dimension $D$ to create a feature vector, applying max-pooling to summarize the mask features, which we term as a ghost masking embedding. For a mask vector $\bm{m}^{t}\in\mathbb{R}^{N\times 1}$ of $N$ agents at time step $t$, the ghost masking embedding is computed as follows:
$\bm{f}^{t}_{gho}=\max_{i=1}^{N}\left(\text{repeat}(\bm{m}^{t},D)[i,:]\right),$ (4)
where the $\text{repeat}(\cdot)$ function expands the mask vector into a $N\times D$ matrix by repeating it $D$ times, and max-pooling is applied across rows to derive the ghost masking embedding $\bm{f}^{t}_{gho}\in\mathbb{R}^{1\times D}$. We also explore other permutation-invariant operations like mean or sum, as shown in Section 4.3, which perform effectively in our model.
@@
To enhance the temporal features extraction while reserving missing relationships, we introduce a Bidirectional Temporal Mamba. This adaptation incorporates multiple residual Mamba blocks paired with an innovative Bidirectional Temporal Scaled (BTS) module. Initially, we process the mask $M$ for the entire trajectory by reversing it along the time dimension to produce $\overleftarrow{M}$, which facilitates the model’s learning of temporal missing relationships by utilizing both the original and flipped masks within our BTS module. This process generates the scaling matrix $\overrightarrow{S}$ and its corresponding reversed version $\overleftarrow{S}$. Specifically, for agent $i$ at time step $t$, $s^{t}_{i}$ is computed as follows:
@@
[…]
where $\varphi_{s}(\cdot)$ are projection functions with weights $\mathbf{W}_{s}$. We implement $\varphi_{s}(\cdot)$ using MLPs with the ReLU activation function.
Equation 6 is designed to calculate the distance from the last observation to the current time step, which helps in quantifying the influence of temporal gaps, particularly when dealing with complex missing patterns. The insight is that the influence of a variable that has been missing for a period decreases over time. Therefore, we utilize a negative exponential function and the ReLU to ensure that the influence decays monotonically within a reasonable range between 00 and $1$. In this way, the transformation function in the Mamba model [25] can be revised for our purposes as follows:
@@
where $\varphi_{q}$ is implemented using MLPs. We sample latent variables $\bm{Z}\sim\mathcal{N}(\bm{\mu}_{q},\text{Diag}(\bm{\sigma}_{q}^{2}))$ for trajectory generation. During testing, we sample $\bm{Z}$ from the prior Gaussian distribution $\mathcal{N}(0,\mathbf{I})$.
@@
where $\varphi_{dec}$ is the decoder function implemented using MLPs.
@@ 4.1 Benchmarks and Setup
We curate and benchmark three sports game datasets for this integrated challenging trajectory generation task, Basketball-U, Football-U, and Soccer-U. (1) Basketball-U: We build Basketball-U from NBA dataset [90], with 93,490 training sequences and 11,543 testing sequences. Each sequence consists of trajectories for 1 ball, 5 offensive players, and 5 defensive players. (2) Football-U: Football-U is created from NFL Football Dataset  [^1] https://github.com/nfl-football-ops/Big-Data-Bowl, with 10,762 training sequences and 2,624 testing sequences. Each sequence consists of trajectories for 1 ball, 11 offensive players, and 11 defensive players. (3) Soccer-U: SoccerTrack  [^2] https://github.com/AtomScott/SportsLabKit [63] dataset is used as the base to build Soccer-U. The top-view scenarios are used to extract 9,882 training sequences and 2,448 testing sequences. Each sequence consists of trajectories for 1 ball, 11 offensive players, and 11 defensive players. To cover various input situations, we design different 5 masking strategies. In our datasets, we set the trajectory length to $T=50$, and the number of agents $N$ includes all the players and the ball. Details can be found in Appendix B.
@@
We use the following three categories of methods for comparison. Although some of these models were not originally designed for our trajectory generation task, we have adapted them to fit our needs. (1) Statistical approach: This includes basic methods such as Mean, Median, and Linear Fit, (2) Vanilla models: These are fundamental networks such as LSTM [33] and Transformer [71], (3) Advanced baselines: Advanced deep learning models such as MAT [90], Naomi [43], INAM [58], SSSD [3], and GC-VRNN [85].

### v2

# Sports-Traj: A Unified Trajectory Generation Model for Multi-Agent Movement in Sports

@@ Abstract
Understanding multi-agent movement is critical across various fields. The conventional approaches typically focus on separate tasks such as trajectory prediction, imputation, or spatial-temporal recovery. Considering the unique formulation and constraint of each task, most existing methods are tailored for only one, limiting the ability to handle multiple tasks simultaneously, which is a common requirement in real-world scenarios. Another limitation is that widely used public datasets mainly focus on pedestrian movements with casual, loosely connected patterns, where interactions between individuals are not always present, especially at a long distance, making them less representative of more structured environments. To overcome these limitations, we propose a Unified Trajectory Generation model, UniTraj, that processes arbitrary trajectories as masked inputs, adaptable to diverse scenarios in the domain of sports games. Specifically, we introduce a Ghost Spatial Masking (GSM) module, embedded within a Transformer encoder, for spatial feature extraction. We further extend recent State Space Models (SSMs), known as the Mamba model, into a Bidirectional Temporal Mamba (BTM) to better capture temporal dependencies. Additionally, we incorporate a Bidirectional Temporal Scaled (BTS) module to thoroughly scan trajectories while preserving temporal missing relationships. Furthermore, we curate and benchmark three practical sports datasets, Basketball-U, Football-U, and Soccer-U, for evaluation. Extensive experiments demonstrate the superior performance of our model. We hope that our work can advance the understanding of human movement in real-world applications, particularly in sports. Our datasets, code, and model weights are available here.
@@ 1 Introduction
Understanding multi-agent movement patterns is invaluable across various domains, including autonomous driving (Codevilla et al., 2019; Hazard et al., 2022; Xu et al., 2024; Yang et al., 2024), video surveillance (Cristani et al., 2013; Coşar et al., 2016), and sports analytics (Tuyls et al., 2021; Wang et al., 2024b). To decipher agent movement, these applications rely on tasks such as multi-object tracking (Luo et al., 2021), person re-identification (Ye et al., 2021), trajectory modeling (Nagin, 2010), and action recognition (Zhang et al., 2022; Chi et al., 2023). Among these tasks, trajectory modeling is particularly straightforward and effective for understanding the movements. Despite its inherent challenges, such as the complexity of dynamic environments and the subtle agent interactions, significant advancements have been made recently. These advancements are concentrated in three main areas: trajectory prediction, imputation, and spatial-temporal recovery.
Significant developments have been made recently in modeling trajectories, yet most approaches are specialized for a single task. For example, numerous studies (Xu et al., 2020; 2021; 2022b; Rowe et al., 2023; Mao et al., 2023; Jiang et al., 2023; Chen et al., 2023a; Zhou et al., 2023; Gu et al., 2023; Aydemir et al., 2023; Bae et al., 2023; Shi et al., 2023; Chen et al., 2023b; Seff et al., 2023; Park et al., 2023; Xu et al., 2023a; Maeda & Ukita, 2023; Park et al., 2024; Xu & Fu, 2024) have focused on pedestrian trajectory prediction, driven by the growing interest in autonomous driving, achieving promising results on public datasets. However, these approaches often struggle to generalize to other trajectory-related tasks, such as trajectory imputation and spatial-temporal recovery. These tasks require exploiting both forward and backward spatial-temporal dependencies, which are not typically
@@
addressed in prediction models. Moreover, early datasets like ETH-UCY (Pellegrini et al., 2009; Lerner et al., 2007) and SDD (Robicquet et al., 2016) primarily focus on pedestrian trajectories in real-world scenarios, such as university campuses and sidewalks. In these settings, pedestrians typically move casually, with limited sparse interactions between individuals, particularly those at longer distances. The most common social interactions observed are group walking or collision avoidance. While some studies (Liu et al., 2019; Chib & Singh, 2025) have tackled the multi-agent trajectory imputation problem, they often overlook the future trajectories of agents. This omission limits their practical utility in complete movement understanding, where predicting future trajectories is crucial for downstream planning rather than merely reconstructing historical trajectories. Although some recent efforts (Xu et al., 2023b; Qin et al., 2023) have attempted to integrate imputation and prediction, they both rely on a multi-task framework, and the masking strategies fall short in handling diverse missing patterns. Given that any situation might happen in real practice, it is crucial to develop a general method that can accommodate various scenarios as shown in Figure 1. This raises two pivotal questions: (1) How can we unify these disparate but relevant tasks into a general framework that accommodates different settings? (2) How can we effectively model these trajectories with varying input formulations?
Prompted by these questions, we revisit these tasks and introduce the Unified Trajectory Generation model (UniTraj), which integrates these tasks into a general scheme. Specifically, we merge different input types into a single unified formulation: an arbitrary incomplete trajectory with a mask that indicates the visibility of each agent at each time step. We uniformly process the inputs of each task as masked trajectories, aiming to generate complete trajectories based on the incomplete ones. To model spatial-temporal dependencies across various trajectory formulations, we introduce a Ghost Spatial Masking (GSM) module embedded within a Transformer-based encoder for spatial feature extraction. Leveraging the notable capability of recent popular State Space Models (SSMs), namely the Mamba model, we adapt and enhance it into a Bidirectional Temporal Mamba encoder for long-term multi-agent trajectory generation. Furthermore, we propose a simple yet effective Bidirectional Temporal Scaled (BTS) module that comprehensively scans trajectories while preserving the integrity of temporal relationships within the sequence. Due to the lack of densely structured trajectory datasets, we benchmark and release three sports datasets: Basketball-U, Football-U, and Soccer-U, to facilitate evaluation. In summary, the contributions of our work are as follows:
@@
  • We curate and benchmark three sports datasets, Basketball-U, Football-U, and Soccer-U, and establish strong baselines for this integrated challenge.
@@ 2.1 Problem Definition
Figure 2: Overall architecture of our UniTraj model. The encoders extract agent features and derive latent variables, while the decoder generates the complete trajectory using the sampled latent variables and agent features.
@@ 2.2 Unified Trajectory Generation Model
where $\varphi_{x}(\cdot)$ and $\varphi_{y}(\cdot)$ are projection functions with weights $\mathbf{W}_{x}\in\mathbb{R}^{8\times D}$ and $\mathbf{W}_{y}\in\mathbb{R}^{8\times D}$, $\odot$ represents element-wise multiplication, and $\oplus$ indicates concatenation. We implemented $\varphi_{x}(\cdot)$ and $\varphi_{y}(\cdot)$ using an MLP. This approach allows us to incorporate location, velocity, visibility, and category information for extracting spatial features for subsequent analyses.
@@
Unlike other sequential modeling tasks, it is crucial to consider dense social interactions, especially in sports games. Existing studies on human interactions predominantly use attention mechanisms like cross-attention and graph attention to capture these dynamics. However, given that we are addressing a unified problem with arbitrary incomplete inputs, it is essential for our model to learn the spatial-temporal missing patterns. While some studies (Zhao et al., 2022; Xu et al., 2023b) utilize edge graphs with attached labels to extract missing features, this approach is resource-intensive and requires the construction of an additional graph layer. In contrast, we propose a novel and efficient Ghost Spatial Masking (GSM) module to abstract and summarize the spatial structures of missing data. This module can be smoothly integrated into the Transformer without complicating the model structure.
The Transformer (Vaswani et al., 2017) was originally proposed to model the temporal dependencies for sequential data, and we adopt the multi-head self-attention design along the spatial dimension. At each time step, we treat each agent’s embedding as a token and input $N$ agent tokens into the Transformer encoder to capture the attentive relationships among the agents’ features. In this way, agent-agent interactions are computed through the multi-head self-attention mechanism. However, it is important to note that at each time step, the missing patterns vary, with some players considered absent. A straightforward approach is to encode these missing features at each time step as well. Given the current mask vector $\bm{m}^{t}\in\mathbb{R}^{N\times 1}$ of $N$ agents at time step $t$, our goal is to generate a masking embedding based on this mask. We propose a simple yet effective method to generate the so-called ghost masking embedding as follows:
$\bm{f}^{t}_{gho}=\max_{i=1}^{N}\left(\text{MLP}\left(\text{repeat}(\bm{m}^{t},D)[i,:]\right)\right),$ (4)
where the $\text{repeat}(\cdot)$ function expands the mask vector into a $N\times D$ matrix by repeating it $D$ times, MLP is a linear layer used to project the mask matrix, and max-pooling is applied across rows to derive the ghost masking embedding $\bm{f}^{t}_{gho}\in\mathbb{R}^{1\times D}$. We also explore alternative operations, such as mean or sum pooling, as described in Section 3.3, which also perform effectively in our model.
@@
To enhance the temporal features extraction while reserving missing relationships, we introduce a Bidirectional Temporal Mamba. This adaptation incorporates multiple residual Mamba blocks paired with an innovative Bidirectional Temporal Scaled (BTS) module. The insight behind this design is that the Mamba blocks are used to extract temporal features, but it is also essential to learn the missing patterns of each trajectory, as these patterns vary across different players. The BTS module is specifically designed to capture these temporal missing patterns, embedding them within the Mamba blocks to ensure more accurate temporal modeling.
Initially, we process the mask $M$ for the entire trajectory by reversing it along the time dimension to produce $\overleftarrow{M}$, which facilitates the model’s learning of temporal missing relationships by utilizing both the original and flipped masks within our BTS module. This process generates the scaling matrix $\overrightarrow{S}$ and its corresponding reversed version $\overleftarrow{S}$. Specifically, for agent $i$ at time step $t$, $s^{t}_{i}$ is computed as follows:
@@
$\overleftrightarrow{F}_{bts}=1/\exp\left(\varphi_{s}(\overleftrightarrow{S};\mathbf{W}_{s})\right),$ (7)
where $\varphi_{s}(\cdot)$ is a projection function with weights $\mathbf{W}_{s}$, $\overleftrightarrow{S}=[\overrightarrow{S},\overleftarrow{S}]$, and $\overleftrightarrow{F}_{bts}=[\overrightarrow{F}_{bts},\overleftarrow{F}_{bts}]$. We implement $\varphi_{s}(\cdot)$ using an MLP with the ReLU activation function.
Equation 6 is designed to calculate the distance from the last observation to the current time step, which helps in quantifying the influence of temporal gaps, particularly when dealing with complex missing patterns. The insight is that the influence of a variable that has been missing for a period decreases over time. Therefore, we utilize a negative exponential function with ReLU to ensure that the influence decays monotonically within a reasonable range between 0 and 1. Accordingly, the transformation function in Mamba model (Gu & Dao, 2023) is revised as follows:
@@
where $\varphi_{q}$ is implemented using an MLP. We sample latent variables $\bm{Z}\sim\mathcal{N}(\bm{\mu}_{q},\text{Diag}(\bm{\sigma}_{q}^{2}))$ for trajectory generation. During testing, we sample $\bm{Z}$ from the prior Gaussian distribution $\mathcal{N}(0,\mathbf{I})$.
@@
where $\varphi_{dec}$ is the decoder function implemented using an MLP.
@@ 3.1 Benchmarks and Setup
We curate and benchmark three sports game datasets for this integrated challenging trajectory generation task, Basketball-U, Football-U, and Soccer-U. (1) Basketball-U: We build Basketball-U from NBA dataset (Zhan et al., 2018), with 93,490 training sequences and 11,543 testing sequences. Each sequence consists of trajectories for 1 ball, 5 offensive players, and 5 defensive players. (2) Football-U: Football-U is created from NFL Football Dataset  [^1] https://github.com/nfl-football-ops/Big-Data-Bowl, with 10,762 training sequences and 2,624 testing sequences. Each sequence consists of trajectories for 1 ball, 11 offensive players, and 11 defensive players. (3) Soccer-U: SoccerTrack  [^2] https://github.com/AtomScott/SportsLabKit (Scott et al., 2022) dataset is used as the base to build Soccer-U. The top-view scenarios are used to extract 9,882 training sequences and 2,448 testing sequences. Each sequence consists of trajectories for 1 ball, 11 offensive players, and 11 defensive players. To cover various input situations, we design different 5 masking strategies. In our datasets, we set the trajectory length to $T=50$, and the number of agents $N$ includes all the players and the ball. Details can be found in Appendix C.
@@
We use the following three categories of methods for comparison. Although some of these models were not originally designed for our trajectory generation task, we have adapted them to fit our needs. (1) Statistical approach: This includes basic methods such as Mean, Median, and Linear Fit, (2) Vanilla models: These are fundamental networks such as LSTM (Hochreiter & Schmidhuber, 1997) and Transformer (Vaswani et al., 2017), (3) Advanced baselines: Advanced deep learning models such as MAT (Zhan et al., 2018), Naomi (Liu et al., 2019), INAM (Qi et al., 2020), SSSD (Alcaraz & Strodthoff, 2022), and GC-VRNN (Xu et al., 2023b).
Evaluation Protocol.
We generate a total of $K=20$ trajectories based on the masked trajectory input and use the minADE20 as one of the evaluation metrics. Furthermore, we include four additional metrics (Zhan et al., 2018; Qi et al., 2020) to comprehensively assess the generation quality. (1) minADE20: Calculate the minimum average displacement error between the generated trajectories and the ground truth across the 20 generated trajectories, (2) Out-of-Boundary (OOB): Measure the percentage of generated locations that fall outside the predefined court boundaries, (3) Step: Calculate the average change in step size across the generated trajectories, (4) Path-L: Calculate the average length of the trajectories for each agent, (5) Path-D: Calculate the maximum difference in trajectory lengths among the agents. Note that except for the metrics minADE20 and OOB, where lower values are better, other metrics typically perform better the closer they are to the ground truth. Implementation details are provided in Appendix D.
3.2 Main Results